OpenStudy (anonymous):

Rewrite the logarithmic equation log4(1/16) = -2 in exponential form.

hartnn (hartnn):

use this \(\Large \log_ab=c \implies b =a^c\) so, \(\large \log_4(1/16) = -2 \implies ????\)

OpenStudy (jdoe0001):

\(\bf log_{\color{red}{ a}}{\color{blue}{ b}}=y\implies {\color{red}{ a}}^y={\color{blue}{ b}} \\ \quad \\ log_{\color{red}{ 4}}\left({\color{blue}{ \frac{1}{16}}}\right)=-2i\implies ?\)